OpenStudy (anonymous):

Solve using the square root method. 9x^2 - 25 = 0 A. {-5/3, 5/3} B. {-3/5, 3/5} C. {-3, 5} D. {-5, 3}

OpenStudy (anonymous):

i think its C.

jimthompson5910 (jim_thompson5910):

\[\Large 9x^2 - 25 = 0\] \[\Large 9x^2 - 25+25 = 0+25\] \[\Large 9x^2 = 25\] \[\Large \frac{9x^2}{9} = \frac{25}{9}\] \[\Large x^2 = \frac{25}{9}\] \[\Large \sqrt{x^2} = \pm\sqrt{\frac{25}{9}}\] \[\Large x = \frac{5}{3} \ \text{or} \ x = -\frac{5}{3}\]